michael is 4 years older than tom. 5 years ago michael was twice toms age how old is tom now
Answer by ankor@dixie-net.com(12695)

Let m = Michael's age now
Let t = toms age now
;
write an equation for each statement:
:
"michael is 4 years older than tom."
m = t + 4
:
"5 years ago michael was twice toms age"
m - 5 = 2(t - 5)
:
m - 5 = 2t - 10
;
Replace m with (t+4); find t
(t+4) - 5 = 2t - 10
;
t - 1 = 2t - 10
:
-1 + 10 = 2t - t
:
t = 9 yrs is Tom's age now
:
Check solution in the statement: (m = 13)
"5 years ago michael was twice toms age"
13 - 5 = 2(9 - 5)